# TOB-SVD: Total-Order Broadcast with Single-Vote Decisions in the Sleepy Model

### v2

@@ 3.2 Graded Agreement and Total-Order Broadcast
Given two logs, $\Lambda$ and $\Lambda^{\prime}$, the notation $\Lambda\preceq\Lambda^{\prime}$ indicates that $\Lambda$ is a prefix of $\Lambda^{\prime}$, implying $\Lambda^{\prime}$ extends $\Lambda$. Two logs are compatible if one acts as a prefix for the other. Conversely, if neither log is a prefix of the other, they are conflicting. Finally, we say that a log $\Lambda$ is an extension of $\Lambda^{\prime}$ if and only if $\Lambda^{\prime}$ is a prefix of $\Lambda$. We assume that any log is an extension of a log $\Lambda_{\mathsf{g}}$ known to any validator. [^8] In blockchain protocols, $\Lambda_{\mathsf{g}}$ typically consists of a log of length $1$.
Graded Agreement. We define a generic Graded Agreement (GA) primitive, with maximum grade $k\geq 1$. In such a primitive, each validator has an input log $\Lambda$, and validators can output logs with grade $0\leq g\leq k$, which we denote with the pair $(\Lambda,g)$.
@@
We require the Graded Agreement primitive to satisfy the following properties.
@@
    Graded Delivery: If an honest validator outputs $(\Lambda,g)$ for $g>0$, any honest validator that participates in the output phase for grade $g-1$ outputs $(\Lambda,g-1)$.
@@
    Validity: If each honest validator awake at time 00 inputs a potentially different extension of a log $\Lambda$, then all honest validators participating in the output phase for a grade $g$ output $(\Lambda,g)$.
@@
Note that, for $g>0$, Consistency already implies Uniqueness. It is only a separate property for outputs of grade 00.
@@
[…]
@@ 3.3 Common notions
Messages. Our protocol defines only one type of message, the $\mathsf{LOG}$ message $\langle\mathsf{LOG},\Lambda\rangle$ where $\Lambda$ is a log. Informally, we say that validator $v_{i}$ sends log $\Lambda$ to mean that $v_{i}$ sends a message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. Similarly, we say that a validator $v_{j}$ receives a log from validator $v_{i}$ to mean that validator $v_{j}$ receives message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$.
@@
As mentioned at the beginning of this section, such a leader election requires us to consider a mildly adaptive corruption model, where the adversary has a delay of time $\Delta$ between scheduling a corruption and executing it. This appears to be necessary also in other protocols which use this strategy. To see why, consider the usual leader election where VRF values are broadcast at time $t$ and a leader is chosen at time $t+\Delta$ based on the highest VRF value so far observed. Between time $t$ and $t+\Delta$, an adaptive adversary can observe the highest VRF value and corrupt its sender, then have it deliver an equivocating proposal only to a subset of the honest validators. This way, some subset of the honest validators only knows of a single proposal with the highest VRF value, and some other subset knows of two such proposals. We cannot then ensure that all honest validators vote [^10] The terms “propose” and “vote” are used here informally to appeal to the reader’s intuition. These terms will be formally defined in the following section. for the same proposal, which is typically required by the Liveness arguments.
@@
For the following, we define a good leader for view $v$ to be a validator in $H_{t_{v}}\setminus B_{t_{v}+\Delta}$ holding the highest VRF value for view $v$ among validators $H_{t_{v}}\cup B_{t_{v}+\Delta}$, i.e., among all validators from which a proposal for view $v$ might be received by time $t_{v}+\Delta$. Note that a good leader always proposes something.
@@
Validator state. At all times, an honest validator keeps only two local variables, $V$ and $E$. First, $V$ associates to a validator $v_{i}$ the pair $V(i)=({\langle\mathsf{LOG},\Lambda\rangle}_{i},t_{i})$ if it has received unique message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$ (from validator $v_{i}$) at time $t_{i}$, or $V(i)=\bot$ if either none or at least two messages $\langle\mathsf{LOG},\Lambda\rangle$ and $\langle\mathsf{LOG},\Lambda^{\prime}\rangle$ with $\Lambda\neq\Lambda^{\prime}$ have been received from $v_{i}$. In other words, $V$ keeps track of non-equivocating $\mathsf{LOG}$ messages together with their reception time. We write $v_{i}\in V$ if and only if $V(i)\not=\bot$, and write $V_{\Lambda}$ for the set containing all extensions of $\Lambda$ recorded in $V$, paired with their sender, i.e., $V_{\Lambda}=\{(\Lambda^{\prime},v_{i}):v_{i}\in V,V(i)=({\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i},*),\Lambda^{\prime}\succeq\Lambda\}$. We say that a validator $v_{i}$ is in $V_{\Lambda}$ to mean that there exists a pair $(*,v_{i})\in V_{\Lambda}$. Similarly, we say that a log $\Lambda^{\prime}$ is in $V_{\Lambda}$ to mean that there exists a pair $(*,\Lambda^{\prime})\in V_{\Lambda}$.
@@
Moreover, $E$ contains a record of equivocators and equivocation evidence, i.e., $E(i)=\bot$ if $v_{i}$ is not known to have equivocated, and otherwise $E(i)=({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$, where $({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$ is equivocation evidence for validator $v_{i}$. As for $V$, we write $v_{i}\in E$ if and only if $E(i)\not=\bot$. A validator can compute from $V$ and $E$ the set $S=\{v_{i}\in\mathcal{V}:v_{i}\in V\lor v_{i}\in E\}$ of all the senders of $\mathsf{LOG}$ messages, i.e., of all validators from which at least one $\mathsf{LOG}$ message has been received.
@@
When we want to emphasize the time $t$ at which we consider these variables, we write $V^{t}$, $E^{t}$, and $S^{t}$. If we also want to emphasize the validator $v_{i}$ whose sets we consider, we write $V^{t,i}$, $E^{t,i}$, and $S^{t,i}$. Note that a validator $v_{i}$ which was awake at time $t$ can access $V^{t,i}$ at any later time $t^{\prime}$ through $V^{t^{\prime},i}$, by simply considering only entries in $V^{t^{\prime},i}$ with reception time $\leq t$.
@@
Message handling. If message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$ is received at time $t$, we have three possibilities on how to handle it. If $v_{i}\not\in V$, i.e., we have not received a $\mathsf{LOG}$ message from $v_{i}$ yet, we add $\Lambda$ to $V$ and forward the message. If $v_{i}\in V$ and $V(i)=({\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i},t)$ for $\Lambda^{\prime}\neq\Lambda$, i.e., we are first learning about an equivocation from $v_{i}$, we record the equivocation by setting $V(i)=\bot$ and $E(i)=({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$. Moreover, we also forward the message, to make sure other validators also learn about the equivocation. Finally, the message is ignored if $v_{i}\in E$, i.e., we already know $v_{i}$ is an equivocator.
4 Technical outline
4.1 Starting point
@@
Momose and Ren [10] introduced the first quorum-based Graded Agreement protocol working in the sleepy model [11], achieving optimal adversarial resilience. Tolerating dynamic participation necessitates doing away with absolute quorums (e.g., 1/2 of the validators), and instead defining them based on the participation level at a given point in time. However, participants can have different perceived participation levels, making agreement challenging. To overcome this, Momose and Ren [10] introduce the novel time-shifted quorum technique which we now summarize. Note that, to limit the notation that we need to introduce to explain such a protocol, we adapt the original protocol by Momose and Ren to work on logs rather than values like in the original formulation as the key concept of time-shifted quorum that we want to illustrate here is not affected by such a change. We let $X_{\Lambda}$ [^11] In the original work by Momose and Ren [10], this is denoted as $\mathcal{E}(b)$, where $b$ represents a value rather than a log. Similarly, the following terminology equivalences apply: $V_{\Lambda}$ corresponds to ¯$\overline{\mathcal{E}}(b)$ in their work, while $S^{{2\Delta}}$ is identified with $\mathcal{M}_{1}$ and $S^{{3\Delta}}$ is identified with $\mathcal{M}_{3}$ by the authors. be the set of all validators from which we have received a message $\langle\mathsf{LOG},\Lambda^{\prime}\rangle$ with $\Lambda^{\prime}\succeq\Lambda$, regardless of whether we have also received any different $\mathsf{LOG}$ message, i.e., regardless of whether they have equivocated. Then, the protocol executed by an honest validator $v_{i}$ upon input $\Lambda$ is as follows.
@@
    At $t=3\Delta$, output $(\Lambda,1)$ if $|V_{\Lambda}^{\Delta}|>|S^{3\Delta}|/2$, and $(\Lambda,0)$ if the number of votes for a log $\Lambda^{\prime}\succeq\Lambda$ is greater than half of all received votes.
The key insight is that, if an honest validator $v_{i}$ outputs $(\Lambda,1)$ by seeing $|V_{\Lambda}^{\Delta,i}|>|S^{3\Delta,i}|/2$, then an honest validator $v_{j}$ awake at time $2\Delta$ would have $X_{\Lambda}^{2\Delta,j}$ and $S^{2\Delta,j}$ such that $|X_{\Lambda}^{2\Delta,j}|\geq|V_{\Lambda}^{\Delta,i}|>|S^{3\Delta,i}|/2\geq|S^{2\Delta,j}|/2$, and thus would vote for $\Lambda$. The first inequality holds because any $\mathsf{LOG}$ message seen by $v_{i}$ at time $\Delta$, and thus counted in $V_{\Lambda}^{\Delta,i}$, is forwarded and received by $v_{j}$ by time $2\Delta$, and so counted in $X_{\Lambda}^{2\Delta,j}$ as well. The converse applies to $|S^{2\Delta,j}|/2$ and $|S^{3\Delta,i}|/2$, i.e., $|S^{2\Delta,j}|/2$ is determined $\Delta$ time before $|S^{3\Delta,i}|/2$, justifying the last inequality. In other words, an honest validator outputting $(\Lambda,1)$ implies that all honest validators awake at time $2\Delta$ vote for $\Lambda$, and thus, given that $\rho\geq 1$, also that all honest validators output $(\Lambda,0)$.
@@
A compromise of this approach is that it prevents this GA primitive from satisfying the Uniqueness property on values with grade 00. The same limitation applies to the variant presented in [9]. Implementing a Total-Order Broadcast based on these primitives introduces significant complexity.
Our Graded Agreement with grades $k=2$. Our first Graded Agreement protocol, detailed in Section 5.1, is a variant of the graded agreement of MR [10], designed to also satisfy Uniqueness for every grade. The protocol then given in Figure 1.
Upon input $\Lambda$, a validator $v_{i}$ runs the following algorithm whenever awake. All validators awake at time $2\Delta$ participate in the output phase for grade 00. A validator participates in the output phase for grade $1$ at time $3\Delta$ if and only if it was awake also at time $\Delta$. At any time, honest validators forward any message received. Up to two different $\mathsf{LOG}$ messages per sender are forwarded upon reception. 1. Input phase, $(t=0)$: Broadcast ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. 2. $(t=\Delta)$: Store $V^{\Delta}$. 3. Output phase for grade 00, $(t=2\Delta)$: If $|V^{2\Delta}_{\Lambda}|>|S^{2\Delta}|/2$: Output $(\Lambda,0)$. 4. Output phase for grade $1$, $(t=3\Delta)$: If awake at time $\Delta$: If $|V^{\Delta}_{\Lambda}\cap V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, Output $(\Lambda,1)$.

### v3

@@ 3.2 Graded Agreement and Total-Order Broadcast
Given two logs, $\Lambda$ and $\Lambda^{\prime}$, the notation $\Lambda\preceq\Lambda^{\prime}$ indicates that $\Lambda$ is a prefix of $\Lambda^{\prime}$. Two logs are compatible if one acts as a prefix for the other. Conversely, if neither log is a prefix of the other, they are conflicting. Finally, we say that a log $\Lambda^{\prime}$ is an extension of $\Lambda$ if and only if $\Lambda$ is a prefix of $\Lambda^{\prime}$. We assume that any log is an extension of a log $\Lambda_{\mathsf{g}}$ known to any validator. [^11] In blockchain protocols, $\Lambda_{\mathsf{g}}$ typically consists of a log of length $1$.
Graded Agreement. We define a generic Graded Agreement (GA) primitive, with $k\geq 1$ grades. In such a primitive, each validator has an input log $\Lambda$, and validators can output logs with grade $0\leq g<k$, which we denote with the pair $(\Lambda,g)$.
@@
Keeping in mind that honest validators can output multiple logs with the same grade, we require the Graded Agreement primitive to satisfy the following properties.
@@
    Graded Delivery: If an honest validator outputs $(\Lambda,g)$ for $g>0$, any honest validator that participates in the output phase for grade g1g-1 outputs $(\Lambda,g-1)$.
@@
    Validity: If each honest validator awake at time 0 inputs a potentially different extension of a log $\Lambda$, then all honest validators participating in the output phase for a grade $g$ output $(\Lambda,g)$.
@@
Note that, for $g>0$, Consistency already implies Uniqueness. It is only a separate property for outputs of grade 0.
@@
    Liveness: For every valid transaction $\mathit{tx}$ in the pool of transactions, there exists a time $t$ such that all honest validators awake for sufficiently long [^12] The duration “sufficiently long” varies depending on the protocol. after $t$ deliver a log $\Lambda$ that includes (a block that includes) transaction $\mathit{tx}$.
@@ 3.3 Common notions
Messages. Our protocol defines only one type of message, the $\mathsf{LOG}$ message $\langle\mathsf{LOG},\Lambda\rangle$ where $\Lambda$ is a log. Informally, we say that validator $v_{i}$ sends log $\Lambda$ to mean that $v_{i}$ sends a message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. Similarly, we say that a validator $v_{j}$ receives a log from validator $v_{i}$ to mean that validator $v_{j}$ receives the message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$.
@@
As mentioned at the beginning of this section, such a leader election requires us to consider a mildly adaptive corruption model, where the adversary has a delay of time $\Delta$ between scheduling a corruption and executing it. This appears to be necessary also in other protocols which use this strategy [4, 5, 6]. To see why, consider the usual leader election where VRF values are broadcast at time $t$ and a leader is chosen at time $t+\Delta$ based on the highest VRF value so far observed. Between time $t$ and $t+\Delta$, an adaptive adversary can observe the highest VRF value and corrupt its sender, then have it deliver an equivocating proposal only to a subset of the honest validators. This way, some subset of the honest validators only knows of a single proposal with the highest VRF value, and some other subset knows of two such proposals. We cannot then ensure that all honest validators vote [^13] The terms “propose” and “vote” are used here informally to appeal to the reader’s intuition. These terms will be formally defined in the following section. for the same proposal, which is typically required by the Liveness arguments.
@@
For the following, we define a good leader for view $v$ starting at time $t_{v}$ to be a validator in $H_{t_{v}}\setminus B_{t_{v}+\Delta}$ holding the highest VRF value for view $v$ among validators $H_{t_{v}}\cup B_{t_{v}+\Delta}$, i.e., among all validators from which a proposal for view $v$ might be received by time $t_{v}+\Delta$. Note that a good leader always proposes something.
@@
Validator state. At all times, an honest validator keeps only two local variables, $V$ and $E$. First, $V$ associates to a validator $v_{i}$ the log $V(i)={\langle\mathsf{LOG},\Lambda\rangle}_{i}$ if it has received an unique message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$ (from validator $v_{i}$), or $V(i)=\bot$ if either none or at least two messages $\langle\mathsf{LOG},\Lambda\rangle$ and $\langle\mathsf{LOG},\Lambda^{\prime}\rangle$ with $\Lambda\neq\Lambda^{\prime}$ have been received from $v_{i}$. In other words, $V$ keeps track of non-equivocating $\mathsf{LOG}$ messages. We write $v_{i}\in V$ if and only if $V(i)\not$$\bot$, and write $V_{\Lambda}$ for the set containing all extensions of $\Lambda$ recorded in $V$, paired with their sender, i.e., $V_{\Lambda}=\{(\Lambda^{\prime},v_{i}):v_{i}\in V,V(i)={\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i},\Lambda^{\prime}\succeq\Lambda\}$. We say that a validator $v_{i}$ is in $V_{\Lambda}$ to mean that there exists a pair $(*,v_{i})\in V_{\Lambda}$. Similarly, we say that a log $\Lambda^{\prime}$ is in $V_{\Lambda}$ to mean that there exists a pair $(*,\Lambda^{\prime})\in V_{\Lambda}$.
@@
Moreover, $E$ contains a record of equivocators and equivocation evidence, i.e., $E(i)=\bot$ if $v_{i}$ is not known to have equivocated, and otherwise $E(i)=({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$, where $({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$ is equivocation evidence for validator $v_{i}$. As for $V$, we write $v_{i}\in E$ if and only if $E(i)\not$$\bot$. A validator can compute from $V$ and $E$ the set $S=\{v_{i}\in\mathcal{V}:v_{i}\in V\lor v_{i}\in E\}$ of all the senders of $\mathsf{LOG}$ messages, i.e., of all validators from which at least one $\mathsf{LOG}$ message has been received.
@@
When we want to emphasize the time $t$ at which we consider these variables, we write $V^{t}$, $E^{t}$, and $S^{t}$. If we also want to emphasize the validator $v_{i}$ whose sets we consider, we write $V^{t,i}$, $E^{t,i}$, and $S^{t,i}$.
@@
Message handling. If message ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$ is received at time $t$, we have three possibilities on how to handle it. If $v_{i}\not\in V$, i.e., we have not received a $\mathsf{LOG}$ message from $v_{i}$ yet, we add $\Lambda$ to $V$ and forward the message. If $v_{i}\in V$ and $V(i)={\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i}$ for $\Lambda^{\prime}\neq\Lambda$, i.e., we are first learning about an equivocation from $v_{i}$, we record the equivocation by setting $V(i)=\bot$ and $E(i)=({\langle\mathsf{LOG},\Lambda\rangle}_{i},{\langle\mathsf{LOG},\Lambda^{\prime}\rangle}_{i})$. Moreover, we also forward the message, to make sure other validators also learn about the equivocation. Finally, ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$ is ignored if $v_{i}\in E$, i.e., we already know $v_{i}$ is an equivocator.
4 Background
@@
Momose and Ren [12] introduced the first quorum-based Graded Agreement protocol working in the sleepy model [14], achieving optimal adversarial resilience. Tolerating dynamic participation necessitates doing away with absolute quorums (e.g., 1/2 of the validators), and instead defining them based on the participation level at a given point in time. However, validators can have different perceived participation levels, making agreement challenging. To overcome this, Momose and Ren [12] introduce the novel time-shifted quorum technique which we now summarize. Note that, to limit the notation that we need to introduce to explain such a protocol, we adapt the original protocol by Momose and Ren to work on logs rather than values like in the original formulation as the key concept of time-shifted quorum that we want to illustrate here is not affected by such a change. We let $X_{\Lambda}$ [^14] In the original work by Momose and Ren [12], this is denoted as $\mathcal{E}(b)$, where $b$ represents a value rather than a log. Similarly, the following terminology equivalences apply: $V_{\Lambda}$ corresponds to $\overline{\mathcal{E}}(b)$ in their work, while $S^{{2\Delta}}$ is identified with $\mathcal{M}_{1}$ and $S^{{3\Delta}}$ is identified with $\mathcal{M}_{3}$ in [12]. be the set of all validators from which we have received a message $\langle\mathsf{LOG},\Lambda^{\prime}\rangle$ with $\Lambda^{\prime}\succeq\Lambda$, regardless of whether we have also received any different $\mathsf{LOG}$ message, i.e., regardless of whether they have equivocated. At any time, honest validators forward any message received. Then, the protocol executed by an honest validator $v_{i}$ upon inputting $\Lambda$ is as follows.
@@
    At $t=3\Delta$, output $(\Lambda,1)$ if $|V_{\Lambda}^{\Delta}|>|S^{3\Delta}|/2$, and $(\Lambda,0)$ if the number of $\mathsf{VOTE}$ messages for a log $\Lambda^{\prime}\succeq\Lambda$ is greater than half of all received $\mathsf{VOTE}$ messages.
The key insight is that, if an honest validator $v_{i}$ outputs $(\Lambda,1)$ by seeing $|V_{\Lambda}^{\Delta,i}|>|S^{3\Delta,i}|/2$, then an honest validator $v_{j}$ awake at time $2\Delta$ would have $X_{\Lambda}^{2\Delta,j}$ and $S^{2\Delta,j}$ such that $|X_{\Lambda}^{2\Delta,j}|\geq|V_{\Lambda}^{\Delta,i}|>|S^{3\Delta,i}|/2\geq|S^{2\Delta,j}|/2$, and thus would vote for $\Lambda$. The first inequality holds because any $\mathsf{LOG}$ message seen by $v_{i}$ at time $\Delta$, and thus counted in $V_{\Lambda}^{\Delta,i}$, is forwarded and received by $v_{j}$ by time $2\Delta$, and so counted in $X_{\Lambda}^{2\Delta,j}$ as well. The converse applies to $|S^{2\Delta,j}|/2$ and $|S^{3\Delta,i}|/2$, i.e., $|S^{2\Delta,j}|/2$ is determined $\Delta$ time before $|S^{3\Delta,i}|/2$, justifying the last inequality. In other words, an honest validator outputting $(\Lambda,1)$ implies that all honest validators awake at time $2\Delta$ vote for $\Lambda$, and thus, given that $\rho\leq\frac{1}{2}$, also that all honest validators output $(\Lambda,0)$.
@@
A compromise of this approach is that it prevents this GA primitive from satisfying the Uniqueness property on values with grade 0. The same limitation applies to the variant presented in [11]. Implementing a Total-Order Broadcast based on these primitives introduces significant complexity.
5 TOB-SVD
We begin this section by presenting the foundational building block of our total-order broadcast protocol. A detailed analysis of our results can be found in Section 6.
5.1 Graded Agreement with $k=2$ grades.
Our first Graded Agreement protocol is a variant of the graded agreement of MR [12] (Section 4), designed to also satisfy Uniqueness for every grade. The protocol is given in Figure 1. It lasts $3\Delta$ time, and requires that the number of Byzantine validators is less than half the number of active validators. Therefore, it specifically works in the $(3\Delta,0,\frac{1}{2})$-sleepy model.
Upon input $\Lambda$, a validator $v_{i}$ runs the following algorithm whenever awake. All validators awake at time $2\Delta$ participate in the output phase for grade 0. A validator participates in the output phase for grade $1$ at time $3\Delta$ if and only if it was awake also at time $\Delta$. At any time, honest validators forward any message received. Up to two different $\mathsf{LOG}$ messages per sender are forwarded upon reception. 1. Input phase, $(t=0)$: Broadcast ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. 2. $(t=\Delta)$: Store $V^{\Delta}$. 3. Output phase for grade 0, $(t=2\Delta)$: If $|V^{2\Delta}_{\Lambda}|>|S^{2\Delta}|/2$: Output $(\Lambda,0)$. 4. Output phase for grade $1$, $(t=3\Delta)$: If awake at time $\Delta$: If $|V^{\Delta}_{\Lambda}\cap V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, Output $(\Lambda,1)$.